OpenStudy (anonymous):

Hi, can someone help me with an Algebra 1 question? I really need the help.

OpenStudy (readergirl12):

What is it I am only 12 so dont expect much xD

OpenStudy (anonymous):

I hope you know Algebra. xD @readergirl12

OpenStudy (anonymous):

Iris has been studying an invasive population of snails. This particular snail has no local predators, so the population grows wildly. She has observed that the population follows an exponential rate of growth for fifteen years. 1. Create your own exponential function, f(x), which models the snail population. You will need to identify the principal population of the snails and the rate of growth each year. Explain to Iris how your function shows the principal population and the rate of growth, in complete sentences.

OpenStudy (readergirl12):

Um

OpenStudy (readergirl12):

Which algebra is this...1?

OpenStudy (anonymous):

The course is 'Algebra 1'.

OpenStudy (anonymous):

@readergirl12

OpenStudy (readergirl12):

K Imma have to get my dad So you can the answer Are you okay with waiting?

OpenStudy (readergirl12):

@trenafuoco

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

@readergirl12 Just make it quick though and thank you.

OpenStudy (readergirl12):

Matter of fact I can get some one else

OpenStudy (readergirl12):

@ganshie8

OpenStudy (readergirl12):

Fail xD

OpenStudy (anonymous):

Haha xD

OpenStudy (readergirl12):

@iGreen

OpenStudy (readergirl12):

@aaronq

OpenStudy (anonymous):

I hope they are all online.

OpenStudy (anonymous):

@tinybookworm

OpenStudy (readergirl12):

They are supposed to be good at math..I dont wanna get you the wrong answer cause i dont know it

OpenStudy (anonymous):

Its okay; is your Dad around or no?

OpenStudy (readergirl12):

He's looking in his book but thanks gonna take too long

OpenStudy (readergirl12):

Thats

OpenStudy (anonymous):

Oh.

OpenStudy (readergirl12):

www.purplemath.com/modules/expofcns.htm

OpenStudy (anonymous):

Hmmm... I don't understand this.

OpenStudy (readergirl12):

Me either xc

OpenStudy (readergirl12):

Do you know what a exponential function is?

OpenStudy (readergirl12):

www.mathsisfun.com/sets/function-exponential.html

OpenStudy (anonymous):

Uhm I have an idea but it may be wrong.

OpenStudy (readergirl12):

What is it?

OpenStudy (anonymous):

Please tell the idea you have. I've been stuck on it for a while. @tinybookworm

OpenStudy (anonymous):

@readergirl12 I do know what it is. its f(x)=a^x

OpenStudy (anonymous):

Let say the population of snails is P.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

Then the rate of growth is r.

OpenStudy (anonymous):

Then x is the numbers of years. G and r are constant. x is the variable

OpenStudy (anonymous):

We will have the function \(y = G.r^x\)

OpenStudy (anonymous):

Yeah. I kinda see where you are going with this.

OpenStudy (anonymous):

Sorry I mean \(P\) instead of \(G\).

OpenStudy (anonymous):

Okay. :)

OpenStudy (anonymous):

For example, the original population is 1000. And the rate of growth is 1.5

OpenStudy (anonymous):

So the population will be \(1000 \times 1.5=1500\) after the first year.

OpenStudy (anonymous):

Okay. :)

OpenStudy (anonymous):

After the second year, the population will be \(1500\times 1.5 =1000\times 1.5^2=2250\)

OpenStudy (anonymous):

Okay, is that it for this question? I have three more that needs to be answered.

OpenStudy (anonymous):

Okay. I will do my best to help you

OpenStudy (readergirl12):

Did you get it right?

OpenStudy (anonymous):

@readergirl12 Yeah. :)

OpenStudy (readergirl12):

YAY WE DID IT

OpenStudy (anonymous):

2. A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population.

OpenStudy (anonymous):

Yeah, and there is two more after this one.

OpenStudy (anonymous):

Is \(g(x)=200(1.03)^{2x}\)?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

As you know, \((1.03)^{2x}\) can be written as \((1.03^x)^2\) or \((1.03^2)^x\). Can you see how to write the function with only x as the exponent?

OpenStudy (anonymous):

Yes. :)

OpenStudy (anonymous):

\((1.03^2)^x=1.0609^x\). So the function can also be written as \(g(x)= 200\times 1.0609^x\)

OpenStudy (anonymous):

Thank you, is that it for this question?

OpenStudy (anonymous):

@tinybookworm

OpenStudy (anonymous):

Sorry. I can't solve the second part: "Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population."

OpenStudy (anonymous):

Oh Okay, thats fine. I can figure that part out. @tinybookworm

OpenStudy (anonymous):

Let me post the third question.

OpenStudy (anonymous):

3. Iris wants to graph the invasive snail population to show the city council. Justify what the appropriate domain and range would be for the function f(x), what the y-intercept would be, and if the function is increasing or decreasing.

OpenStudy (anonymous):

I think the domain will be \([0,+\infty\)) and the range will be \([P,+\infty)\), with \(P\) is the original population

OpenStudy (anonymous):

Okay. :) And that will be it for this question?

OpenStudy (anonymous):

Oh, the y-intercept will be (0,P) and the function is increasing

OpenStudy (anonymous):

Okay :) Now for the fourth question.

OpenStudy (anonymous):

4. In five years, a garden festival plans on using the park where Iris has been studying the invasive snails. Explain to the garden festival committee how to find the average rate of change for the snail population between years 2 and 5. Describe what this average rate of change represents. @tinybookworm

OpenStudy (anonymous):

I think to find the average change, we can calculate the difference in population between year 5 and year 2, then divide by 3

OpenStudy (anonymous):

You mean the totals from years 2 and 5 that we had earlier? Subtract them, and divide by 3?

OpenStudy (anonymous):

Yes I think

OpenStudy (anonymous):

Okay. Is it that it on that problem?

OpenStudy (anonymous):

Sorry. But I don't know what the rate of change represents.

OpenStudy (anonymous):

Its fine. I'll figure it out. Oh crap I forgot, there is three more questions. Are gonna still stick around for a bit longer for more questions? :) @tinybookworm

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Alright, give me a moment.

OpenStudy (anonymous):

2. Create an exponential growth function, f(x), to model a population of frogs that is growing every year. Identify the principal amount, the growth rate, and the appropriate domain and range for your function. Explain how these key features would affect the graph of f(x).

OpenStudy (anonymous):

I think the function will be \(f(x)=P.r^x\) too. Same domain, \([0,+\infty)\), and range, \([P,+\infty)\).

OpenStudy (anonymous):

Okay. is that it for this question?

OpenStudy (anonymous):

I think we can explain these like what we did to the first problem

OpenStudy (anonymous):

Okay. Lets move on to the next question.

OpenStudy (anonymous):

Using your function f(x) from question 2, demonstrate and explain how to find the average rate of change between year 3 and 5, and between year 5 and 7. Explain what the average rate of change represents to the frog population.

OpenStudy (anonymous):

Same way. Subtract the population in year 5 by the population in year 3 and then divide by 2

OpenStudy (anonymous):

Okay, thank you! :D

OpenStudy (anonymous):

You are very welcome